OpenStudy (anonymous):

Verify the identity [(4sin theta*cos theta)*(1-2sin^2 theta)=sin(4 theta)

OpenStudy (turingtest):

\[4\sin\theta\cos\theta(1-2\sin^2\theta)=\sin(4\theta)\]right?

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (turingtest):

this will undoubtedly help http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

Okay. Thank you.

OpenStudy (turingtest):

look at the double angle formulas. applying two of them to the left side will prove the statement

OpenStudy (anonymous):

Okay, thank you for helping me.